OpenStudy (anonymous):

Which value for x makes the sentence true? 4x - 1 = 23 x = 6 x = 28 x = 24 x = 4

pooja195 (pooja195):

add 1 to 23 then divide

pooja195 (pooja195):

4x/24=?

OpenStudy (black_ninja123):

4*6=24 24-1=23 x=6

OpenStudy (anonymous):

thanks lol

OpenStudy (lizz123):

4x - 1 = 23 Since there is the subtraction you have to do the opposite which is addition so iyou have to add by 1 on both sides. 4x - 1 = 23 4x + 1 = + 1 4x _________ = _________ 4x = 24 You have to divde by 4 both sides. 4x = 24 ____________ ___________ 4 4 x divide it by 4 is cancelled and 24 divided by 4 equals 6. Therefore, x=6.